OpenStudy (anonymous):

Given: y = 3x - 4. What is the x-intercept?

OpenStudy (anonymous):

To find the x-intercept, make the y=0 and solve for x. y=3x-4 0=3x-4 4=3x x=4/3 x-intercept: 4/3 Written as a point, it would be (4/3, 0)

OpenStudy (anonymous):

Oh so you turn the Y into a 0.

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

Thanks for helping see how it's done